The cost of debt Gronseth Drywall Systems, Inc, is in discussions with its investment bankers r informed the firm that different matunities will carry diferent coupon rates and sell at different prices. The firm must choose among several alternatives. In each case the on s wil have a S1 000 par value and tation costs will be $40 per bond The company s taxed at 24% Use the approximation formula to calculate the aff tax cost of financing with the following alternative. (Click on the icon located on the top-rnight cormer of the data table below in order to copy its contents into a spreadsheet ) coupon rate 12% Time to maturity Premium or discount - The after-tax cost of financing using the approximation formula is Round to two decimal places ) Type here
media%2F296%2F29687362-89ee-41e1-9b5c-26
Answer #1
1) Approximate before tax cost of debt = [C+(F-P)/n]/(F+P)/2
Where
C = Coupon payment
F = Face value
P = Price
n = Years to maturity
Substituting values we have,
Approximate before tax cost of debt = ((120+(1000-1000+300+40)/14))/((1000+660)/2)) = 17.38%
After tax cost of debt = 17.38%*(1-24%) = 13.21%
2) After tax cost of debt = 9%*(1-35%) = 5.85%
WACC = 11%*40%+12%*20%+5.85%*40% = 9.14%